# Anomalous bulk-edge correspondence of nonlinear Rice-Mele model

### v1

@@ Abstract
Bulk-edge correspondence (BEC) constitutes a fundamental concept within the domain of topological physics, elucidating the profound interplay between the topological invariants that characterize the bulk states and the emergent edge states. A recent highlight along this research line consists of establishing BEC under the eigenvalue’s nonlinearity in a linear Hamiltonian by introducing auxiliary eigenvalues [ T. Isobe et al., Phys. Rev. Lett. 132, 126601 (2024)]. The purpose of this work aims to extend Isobe’s analysis to uncover BEC of eigenvalue’s nonlinearity in intrinsic nonlinear Hamiltonians. To achieve this, we numerically solve the nonlinear Rice-Mele (RM) model and identify two distinct types of nonlinear eigenvalues: the intrinsically nonlinear eigenvalues and the eigenvalue’s nonlinearity introduced through the incorporation of auxiliary eigenvalues. Furthermore, we establish a novel form of BEC based on these auxiliary nonlinear eigenvalues, which we term the anomalous BEC of a nonlinear physical system. The concept of the anomalous BEC defined herein provides a novel perspective on the intricate interplay between topology and nonlinearity in the context of BEC.
@@ I Introduction
The bulk-edge correspondence (BEC) stands as a cornerstone principle in topological physics [1, 2, 3], positing that the topological invariants of bulk states, such as the Chern number, determine the number and nature of edge states. There is currently a surge of interest in exploring the BEC across a diverse spectrum of physical systems [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26], including interdisciplinary areas like meteorological systems [27]. The fascination with the BEC stems from two primary drivers. Firstly, within topological physics [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39], the BEC is a hallmark feature of topological materials [40, 41], encompassing topological insulators, topological semimetals, and topological superconductors. It serves as a pivotal method for identifying and characterizing these materials. Secondly, the BEC not only sheds light on the intrinsic connection between bulk and edge states but also lays the theoretical groundwork for the application of topological materials. For example, in fields such as quantum computing [42, 43, 44] and spintronics [45], harnessing the edge states of topological materials facilitates low-loss and highly stable electron and spin transport. To date, the BEC has predominantly been explored within the realm of linear quantum systems [46, 47].
@@
Extending our understanding beyond linear quantum systems, the intricate interplay between topology and many-body effects [48, 49, 50, 51, 52, 53] in quantum systems gives rise to a rich tapestry of topological many-body phenomena. This interplay has been the subject of intensive investigation in both fermionic [54, 55, 56, 57, 58] and bosonic systems [59, 60, 61, 62, 63]. These efforts are fueled by the rapid advancements in programmable materials and the increasing integration of topological principles into practical applications. A natural and compelling extension of this research is to delve into how the interplay between nonlinearity (or interactions more broadly) and topology influences the behavior of BEC. This line of inquiry promises to uncover new insights into the fundamental properties of BEC and may pave the way for the discovery of novel topological states in these systems.
@@
In the recent work [64], the authors have ventured into the realm of eigenvalue nonlinearity [64, 65, 66, 67, 68, 69, 70] within the context of linear Hamiltonians. Their exploration offers a fresh perspective on the BEC in the nonlinear quantum domain. A notable finding is that, when the nonlinearity is weak but not negligible, the topological edge states of the auxiliary eigenstates are inherited as physical edge states. This discovery underscores the deep connection between the auxiliary eigenvalues and physical systems. The work carries significant physical implications. It suggests that the method of introducing auxiliary eigenvalues can be generalized to systems belonging to different symmetry classes and dimensions, opening up new avenues for understanding and exploring the BEC in a wider range of physical systems.
@@
Despite the groundbreaking insights offered by the aforementioned study, its focus is restricted to eigenvalue’s nonlinearity within the framework of linear Hamiltonians. This naturally prompts the inquiry: What is the nature of the BEC in the realm of nonlinear Hamiltonians [71, 72] that exhibit nonlinear eigenvalues? This question stands as a fertile ground for exploration, as it holds the promise of uncovering novel mechanisms and phenomena within the topological landscape of nonlinear systems.
@@
In this work, we shift our focus to the nonlinearity of the eigenvalues within nonlinear Rice-Mele (RM) Hamiltonians, offering a distinct perspective compared to Ref. [64]. Our investigation encompasses not only the eigenvalue spectrum but also the BEC. Utilizing numerical methods, we delve into how nonlinearity influences the stability of edge states, adopting an approach similar to Ref. [73]. To achieve this, we numerically solve the nonlinear RM model and identify two distinct types of nonlinear eigenvalues: the intrinsically nonlinear eigenvalues of nonlinear RM model and the auxiliary nonlinear eigenvalues introduced through the concept of auxiliary eigenvalues. Furthermore, we establish a novel form of BEC based on these auxiliary nonlinear eigenvalues, which we term the ‘anomalous BEC’ of a nonlinear physical system. The notion of ‘anomalous’ introduced here provides fresh insights into the interplay between topology and nonlinearity in the context of BEC.
@@
Figure 1: Nonlinear eigenvalue of nonlinear RM Hamiltonian defined in the second line of Table 1 and nonlinear Thouless pumping of soltion. (a1)-(d1): Band structures of the nonlinear eigenvalues for the nonlinear RM Hamiltonian (1), (a2)-(d2): The expected position of the soliton as a function of time over a single period. The parameters are fixed as $J=1$, $\delta=0.5$, $\Delta=1$, $T=2000\pi$ and $\omega=10^{-3}$. (a1)-(d1) [(a2)-(d2)]: The interaction strengths are set to $g=0$, $g=1$, $g=5$ and $g=10$, respectively.
@@
The organizational structure of this paper is outlined as follows: In Sec. II, we provide a thorough introduction to the RM Hamiltonian. In Sec. III, four different types of eigenvalue problems are defined. In particular, directly inspiring by Ref. [64], we give the definition of the anomalous eigenvalue’s nonlinearity for the nonlinear RM Hamiltonian and establish a novel form of BEC based on these auxiliary nonlinear eigenvalues. Sec. IV investigate the effects of nonlinear intensity and model parameter on the BEC of nonlinear RM model, elucidating how these factors affect the system’s behavior. Finally, in Sec. V, we summarize the main results of our investigation.
@@ II Nonlinear Rice-Mele Model
In this work, we are interested in a one-dimensional interacting bosonic chain consisting of $N$ dimer units [1]. At the mean-field level, our model system can be well described by the nonlinear RM Hamiltonian [76, 77, 75, 78, 79, 80, 81, 82, 50, 83, 84, 85, 86, 87] as follows
$H_{\text{non}}$ $=$ $\sum_{n=1}^{2N}\Bigg{\{}-\Big{[}J+(-1)^{n}\delta\sin\left(\omega t\right)\Big{]}(\Psi_{n}^{*}\Psi_{n+1}+\mathit{h.c.})$ (1)
In Hamiltonian (1), the $[\Psi_{n},\Psi_{n+1}]$ denotes the wave function of the dimer located at site $n$ [two atoms in one unit cell]. The $J$ is the uniform hopping amplitude, the $\delta\sin(\omega t)$ is the time-dependent dimerization order, and therefore the terms $[-J-\delta\sin(\omega t)]$ and $[-J+\delta\sin(\omega t)]$ represent the intracell and intercell coupling, respectively. The $\Delta\cos(\omega t)$ labels the time-dependent staggered sublattice potential, and the $g$ represents the strength of the focusing Kerr-like nonlinearity [75]. Furthermore, the $\omega$ in Hamiltonian (1), representing the modulation frequency, is be sufficiently small to ensure adiabatic evolution. Finally, we remark that Hamiltonian (1) is of immediate relevance in the context of recent experiments [50] for investigating interaction-induced Thouless pumping in a dynamical superlattice.
The purpose and emphasis of this work is to extend theoretical framework about the eigenvalue’s nonlinearity and the corresponding BEC of linear Hamiltonian in Ref. [64] to the counterpart of nonlinear cases, i. e. the anomalous eigenvalue’s nonlinearity and the corresponding anomalous BEC based on the nonlinear RM Hamiltonian (1). In this end, we proceed to obtain equations of motion of nonlinear RM Hamiltonian (1) by variation of the Hamiltonian (1) as $i\partial\Psi_{j}/\partial t=\delta H/\delta\Psi^{*}_{j}$ which are described by the following set of nonlinear Schrödinger equations for $j=0,1,...N-1$,
$i\frac{\partial\Psi_{2j}}{\partial t}$ $=$ $-\left(J+\delta\sin\omega t\right)\Psi_{2j+1}-\left(J+\delta\sin\omega t\right)\Psi_{2j-1}-\left[\Delta\cos\omega t+g\left|\Psi_{2j}\right|^{2}\right]\Psi_{2j},$ (2)
$i\frac{\partial\Psi_{2j+1}}{\partial t}$ $=$ $-\left(J+\delta\sin\omega t\right)\Psi_{2j}-\left(J+\delta\sin\omega t\right)\Psi_{2j+2}+\left[\Delta\cos\omega t-g\left|\Psi_{2j+1}\right|^{2}\right]\Psi_{2j+1}.$ (3)
Nonlinear eigenvalue of the nonlinear Hamiltonian defined as $\hat{H}_{\text{non}}\Psi=\mu\Psi$ can be obtained by plugging $\Psi\rightarrow\Psi e^{i\mu t/\hbar}$ into Eqs. (2) and (3). We note that Eqs. (2) and (3) constitute a class of time-dependent Gross-Pitaevskii equations, or alternatively, discrete nonlinear Schrödinger equations. These equations provide a powerful framework for describing the dynamical behavior of Bose-Einstein condensates within an ultracold quantum gas [50], as well as the propagation characteristics of pulsed light through arrays of waveguides [75].
@@ III.1 Four kinds of eigenvalues of linear and nonlinear Hamiltonian
In Sec. II, we presented the nonlinear RM Hamiltonian (1) along with the associated discrete nonlinear Schrödinger equations (2) and (3). The objective of Sec. III is to undertake a thorough examination of the linear and nonlinear eigenvalue problems pertinent to both the linear and nonlinear Hamiltonians of Eq. (1), drawing direct inspiration from Ref. [64]. Table 1 offers a comprehensive overview of the definitions of these problems, serving as a crucial reference for understanding the progression from linear to nonlinear eigenvalue problems within Hamiltonian systems.
@@
We provide an outline of the four distinct types of eigenvalue problems, which are summarized in Table 1, as follows:
@@
(i) Eigenvalue of linear Hamiltonian ($\hat{H}_{\text{lin}}\Psi=E\Psi$) as summarized in the first line of Table 1: The nonlinear RM Hamiltonian (1) simplifies to a linear form when the nonlinearity parameter $g$ is set to zero. In this case, the entire parameter space is spanned by $\delta$ and $\Delta$. The BEC of the linear RM Hamiltonian exhibits bulk properties along with two edge states, as depicted in Fig. 1(a1). The topological invariant relevant for charge pumping is the Berry phase of the lowest band, which becomes singular at the origin of the $\delta$-$\Delta$ plane [1]. Specifically, a trajectory that encloses this singularity corresponds to a nonzero Berry phase and pumps two charges to the neighboring unit cell per pump cycle (equivalent to two atoms in one dimer unit). Conversely, if the trajectory does not enclose the singularity, the Berry phase is zero, and consequently, the pumped charge is also zero.
@@
(ii) Nonlinear eigenvalue of the nonlinear Hamiltonian ($\hat{H}_{\text{non}}\Psi=\mu\Psi$) as summarized in the second line of Table 1: This eigenvalue problem is derived by substituting $\Psi\rightarrow\Psi e^{i\mu t/\hbar}$ into Eqs. (2) and (3). The nonlinear RM Hamiltonian (1), characterized by a nonzero nonlinearity parameter $g$, gives rise to a plethora of interaction-induced phenomena that have no counterpart in the linear regime. The synergy between topology and interaction has ushered in new avenues for exploring topological transport and BEC. A quintessential example is the nonlinear Thouless pumping [48]. In this context, nonlinearity plays a pivotal role in quantizing transport through the formation of solitons and spontaneous symmetry-breaking bifurcations (see Figs. 1(b2), (c2), and (d2)). For weak nonlinearity, the motion of the soliton becomes topologically quantized, as illustrated in Figs. 1(b2). Specifically, the BEC of the nonlinear RM Hamiltonian (1) exhibits a quantized displacement by two lattice sites per cycle (see Fig. 1(b2)), which is directly linked to the topology of the underlying band structures (see Fig. 1(b1)). In contrast, no such quantized pumping of solitons occurs when $g=0$, as shown in Fig. 1(a2). As the nonlinearity strength $g$ increases, self-intersecting bands emerge in the intermediate nonlinear regime with the appearance of the loop structure [88] as shown in Fig. 1(c1), leading to a pumping effect that persists but is no longer quantized, as observed in Fig. 1(c2). When the nonlinearity further intensifies into the strong nonlinear regime (see Fig. 1(d1)) and the self-trapped phenomenon is supposed to occur [89], the soliton dynamics are suppressed, resulting in the cessation of the pumping effect, as depicted in Fig. 1(d2).
(iii) Eigenvalue’s nonlinearity in a linear Hamiltonian as summarized in the third line of Table 1: This scenario introduces a novel concept where the eigenvalues exhibit a state-variable-dependent nonlinearity. The eigenvalue equation is given by $\hat{H}_{\text{lin}}\Psi=\omega S(\omega)\Psi$, where $\omega$ is not a simple scalar but a function related to a parameter within the linear Hamiltonian $H_{\text{lin}}$. Remarkably, Ref. [64] has shown that when the nonlinearity is weak but finite, the topological edge states of the auxiliary eigenstates are topologically inherited as physical edge states. This inheritance occurs when the auxiliary eigenvalues are monotonic with respect to the physical eigenvalues. This finding establishes a BEC in systems with eigenvalue nonlinearity, highlighting the profound implications of such nonlinearity on the topological properties of the system.
(iv) Anomalous eigenvalue’s nonlinearity of nonlinear Hamiltonian in the fourth line of Table 1: This scenario extends the concept of eigenvalue nonlinearity from linear Hamiltonians in Ref. [64] to their nonlinear counterparts, which is referred as the anomalous eigenvalue’s nonlinearity of nonlinear Hamiltonian. The definition given by $\hat{H}_{\text{non}}\Psi=\omega S(\omega)\Psi$ represents a generalization where the Hamiltonian is nonlinear. This equation models systems with anharmonic potentials, where energy levels and wave functions are intricately intertwined in a nonlinear fashion, providing a richer framework for understanding quantum states.
@@
Figure 2: Eigenvalue’s nonlinearity of nonlinear RM Hamiltonian defined the last line of Table 1 and nonlinear Thouless pumping of soltion. (a1)-(d1): The auxiliary $\lambda$-spectrum for the nonlinear RM Hamiltonian (1). (a2)-(d2): The anticipated position of the nonlinear excitation of soliton as a function of time over one period. The parameters are fixed at $J=1$, $\delta=0.5$, $\Delta=1$, $T=2000\pi$ and $\omega=10^{-3}$. Specifically, (a1)-(d1) and their corresponding (a2)-(d2) panels represent interaction strengths of $g=0$, $g=1$, $g=5$ and $g=10$, respectively.

### v2

@@ Abstract
Bulk-edge correspondence constitutes a fundamental concept within the domain of topological physics, elucidating the profound interplay between the topological invariants that characterize the bulk states and the emergent edge states. A recent highlight along this research line consists of establishing bulk-edge correspondence under the eigenvalue’s nonlinearity in a linear Hamiltonian by introducing auxiliary eigenvalues [ T. Isobe et al., Phys. Rev. Lett. 132, 126601 (2024)]. The purpose of this work aims to extend Isobe’s analysis to uncover bulkedge correspondence of the eigenvalue’s nonlinearity in intrinsic nonlinear Hamiltonians. To achieve this, we numerically solve the nonlinear Rice-Mele (RM) model and identify two distinct types of nonlinear eigenvalues: the intrinsically nonlinear eigenvalues and the eigenvalue’s nonlinearity introduced through the incorporation of auxiliary eigenvalues. Furthermore, we establish a form of bulk-edge correspondence based on these auxiliary nonlinear eigenvalues, which we term the anomalous bulk-edge correspondence of a nonlinear physical system. The concept of the anomalous bulk-edge correspondence defined herein provides an alternative perspective on the intricate interplay between topology and nonlinearity in the context of bulk-edge correspondence.
@@ I Introduction
The bulk-edge correspondence stands as a cornerstone principle in topological physics [1, 2, 3], positing that the topological invariants of bulk states, such as the Chern number, determine the number and nature of edge states. There is currently a surge of interest in exploring the bulk-edge correspondence across a diverse spectrum of physical systems [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26], including interdisciplinary areas like meteorological systems [27]. The fascination with the bulk-edge correspondence stems from two primary drivers. First, within topological physics [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39], the bulk-edge correspondence is a hallmark feature of topological materials [40, 41], encompassing topological insulators, topological semimetals, and topological superconductors. It serves as a pivotal method for identifying and characterizing these materials. Second, the bulk-edge correspondence not only sheds light on the intrinsic connection between bulk and edge states, but also lays the theoretical groundwork for the application of topological materials. For example, in fields such as quantum computing [42, 43, 44] and spintronics [45], harnessing the edge states of topological materials facilitates low-loss and highly stable electron and spin transport. To date, the bulk-edge correspondence has predominantly been explored within the realm of linear quantum systems [46, 47].
@@
Extending our understanding beyond linear quantum systems, the intricate interplay between topology and many-body effects [48, 49, 50, 51, 52, 53] in quantum systems gives rise to a rich tapestry of topological many-body phenomena. This interplay has been the subject of intensive investigation in both fermionic [54, 55, 56, 57, 58] and bosonic systems [59, 60, 61, 62, 63]. These efforts are fueled by the rapid advancements in programmable materials and the increasing integration of topological principles into practical applications. A natural and compelling extension of this research is to delve into how the interplay between nonlinearity (or interactions more broadly) and topology influences the behavior of bulk-edge correspondence. This line of inquiry promises to uncover additional insights into the fundamental properties of bulk-edge correspondence and may pave the way for the discovery of novel topological states in these systems.
@@
In the recent work [64], the authors have ventured into the realm of eigenvalue nonlinearity [64, 65, 66, 67, 68, 69, 70] within the context of linear Hamiltonians. Their exploration offers a fresh perspective on the bulk-edge correspondence in the nonlinear quantum domain. A notable finding is that, when the nonlinearity is weak but not negligible, the topological edge states of the auxiliary eigenstates are inherited as physical edge states. This discovery underscores the deep connection between the auxiliary eigenvalues and physical systems. The work carries significant physical implications. It suggests that the method of introducing auxiliary eigenvalues can be generalized to systems belonging to different symmetry classes and dimensions, opening up new avenues for understanding and exploring the bulk-edge correspondence in a wider range of physical systems.
@@
Despite the groundbreaking insights offered by the aforementioned study, its focus is restricted to eigenvalue’s nonlinearity within the framework of linear Hamiltonians. This naturally prompts the inquiry: What is the nature of the bulk-edge correspondence in the realm of nonlinear Hamiltonians [71, 72] that exhibit nonlinear eigenvalues? This question stands as a fertile ground for exploration, as it holds the promise of uncovering novel mechanisms and phenomena within the topological landscape of nonlinear systems.
@@
In this work, we shift our focus to the nonlinearity of the eigenvalues within nonlinear Rice-Mele (RM) Hamiltonians, offering a distinct perspective compared to Ref. [64]. Our investigation encompasses not only the eigenvalue spectrum, but also the bulk-edge correspondence. Utilizing numerical methods, we delve into how nonlinearity influences the stability of edge states, adopting an approach similar to Ref. [73]. To achieve this, we numerically solve the nonlinear RM model and identify two distinct types of nonlinear eigenvalues: the intrinsically nonlinear eigenvalues of nonlinear RM model and the auxiliary nonlinear eigenvalues introduced through the concept of auxiliary eigenvalues. Furthermore, we establish a form of bulk-edge correspondence based on these auxiliary nonlinear eigenvalues, which we term the “anomalous bulk-edge correspondence” of a nonlinear physical system. The notion of “anomalous” introduced here provides fresh insights into the interplay between topology and nonlinearity in the context of bulk-edge correspondence.
@@
Figure 1: Nonlinear eigenvalue of nonlinear RM Hamiltonian defined in the second line of Table 1 and nonlinear Thouless pumping of soliton. (a1)-(d1) Band structures of the nonlinear eigenvalues for nonlinear RM Hamiltonian (1). (a2-d2) The expected position of the soliton as a function of time over a single period. The parameters are fixed as $J=1$, $\delta=0.5$, $\Delta=1$, $T=2000\pi$, and $\omega=10^{-3}$. (a1)-(d1)[(a2)-(d2)] The interaction strengths are set to $g=0$, $g=1$, $g=5$, and $g=10$, respectively.
@@
The organizational structure of this paper is outlined as follows. In Sec. II, we provide a thorough introduction to the RM Hamiltonian. In Sec. III, four different types of eigenvalue problems are defined. In particular, directly inspired by the work of the authors of Ref. [64], we give the definition of the anomalous eigenvalue’s nonlinearity for the nonlinear RM Hamiltonian and establish an additional form of bulk-edge correspondence based on these auxiliary nonlinear eigenvalues. Section IV investigate the effects of nonlinear intensity and model parameter on the bulk-edge correspondence of the nonlinear RM model, elucidating how these factors affect the system’s behavior. Finally, in Sec. V, we summarize the main results of our investigation.
@@ II Nonlinear Rice-Mele Model
In this work, we are interested in a one-dimensional interacting bosonic chain consisting of $N$ dimer units [1]. At the mean-field level, our model system can be well described by the nonlinear RM Hamiltonian[74, 75, 76, 77, 78, 79, 80, 81, 50, 82, 83, 84, 85, 86] as follows:
[…]
In Hamiltonian (1), the $[\Psi_{n},\Psi_{n+1}]$ denotes the wave function of the dimer located at site $n$ (two atoms in one unit cell). The $J$ is the uniform hopping amplitude, the $\delta\sin(\omega t)$ is the time-dependent dimerization order, and therefore, the terms $[-J-\delta\sin(\omega t)]$ and $[-J+\delta\sin(\omega t)]$ represent the intracell and intercell coupling, respectively. The $\Delta\cos(\omega t)$ labels the time-dependent staggered sublattice potential, and the $g$ represents the strength of the focusing Kerr-like nonlinearity [76]. Furthermore, the $\omega$ in Hamiltonian (1), representing the modulation frequency, is sufficiently small to ensure adiabatic evolution. Finally, we remark that Hamiltonian (1) is of immediate relevance in the context of recent experiments [50] for investigating interaction-induced Thouless pumping in a dynamical superlattice.
@@
The purpose and emphasis of this work is to extend the theoretical framework about the eigenvalue’s nonlinearity and the corresponding bulk-edge correspondence of the linear Hamiltonian in Ref. [64] to the counterpart of nonlinear cases, i.e., the anomalous eigenvalue’s nonlinearity and the corresponding anomalous bulk-edge correspondence based on the nonlinear RM Hamiltonian (1). In this end, we proceed to obtain equations of motion of the nonlinear RM Hamiltonian (1) by variation of the Hamiltonian (1) as $i\partial\Psi_{j}/\partial t=\delta H/\delta\Psi^{*}_{j}$ which are described by the following set of nonlinear Schrödinger equations for $j=0,1,...N-1$,
$i\frac{\partial\Psi_{2j}}{\partial t}=$displaystyle-- $\left(J+\delta\sin\omega t\right)\Psi_{2j+1}-\left(J-\delta\sin\omega t\right)\Psi_{2j-1}$ (2)
-- $\left[\Delta\cos\omega t+g\left|\Psi_{2j}\right|^{2}\right]\Psi_{2j},$
$i\frac{\partial\Psi_{2j+1}}{\partial t}=$displaystyle-- $\left(J+\delta\sin\omega t\right)\Psi_{2j}-\left(J-\delta\sin\omega t\right)\Psi_{2j+2}$ (3)
$+$ $\left[\Delta\cos\omega t-g\left|\Psi_{2j+1}\right|^{2}\right]\Psi_{2j+1}.$
The nonlinear eigenvalue of the nonlinear Hamiltonian defined as $\hat{H}_{\text{non}}\Psi=\mu\Psi$ can be obtained by plugging $\Psi\rightarrow\Psi e^{i\mu t/\hbar}$ into Eqs. (2) and (3). We note that Eqs. (2) and (3) constitute a class of time-dependent Gross-Pitaevskii equations, or alternatively, discrete nonlinear Schrödinger equations. These equations provide a powerful framework for describing the dynamical behavior of Bose-Einstein condensates within an ultracold quantum gas [50], as well as the propagation characteristics of pulsed light through arrays of waveguides [76].
@@ III.1 Four kinds of eigenvalues of linear and nonlinear Hamiltonian
In Sec. II, we present the nonlinear RM Hamiltonian (1) along with the associated discrete nonlinear Schrödinger equations (2) and (3). The objective of Sec. III is to undertake a thorough examination of the linear and nonlinear eigenvalue problems pertinent to both the linear and nonlinear Hamiltonians of Eq. (1), drawing direct inspiration from Ref. [64]. Table 1 offers a comprehensive overview of the definitions of these problems, serving as a crucial reference for understanding the progression from linear to nonlinear eigenvalue problems within Hamiltonian systems.
@@
We provide an outline of the four distinct types of eigenvalue problems, which are summarized in Table 1, as follows.
@@
(i) The eigenvalue of the linear Hamiltonian ($\hat{H}_{\text{lin}}\Psi=E\Psi$) as summarized in the first line of Table 1: The nonlinear RM Hamiltonian (1) simplifies to a linear form when the nonlinearity parameter $g$ is set to zero. In this case, the entire parameter space is spanned by $\delta$ and $\Delta$. The bulk-edge correspondence of the linear RM Hamiltonian exhibits bulk properties along with two edge states, as depicted in Fig. 1(a1). The topologically invariant relevant for charge pumping is the Berry phase of the lowest band, which becomes singular at the origin of the $\delta$-$\Delta$ plane [1]. Specifically, a trajectory that encloses this singularity corresponds to a nonzero Berry phase and pumps two charges to the neighboring unit cell per pump cycle (equivalent to two atoms in one dimer unit). Conversely, if the trajectory does not enclose the singularity, the Berry phase is zero, and consequently, the pumped charge is also zero.
@@
(ii) Nonlinear eigenvalue of the nonlinear Hamiltonian ($\hat{H}_{\text{non}}\Psi=\mu\Psi$) as summarized in the second line of Table 1: This eigenvalue problem is derived by substituting $\Psi\rightarrow\Psi e^{i\mu t/\hbar}$ into Eqs. (2) and (3). The nonlinear RM Hamiltonian (1), characterized by a nonzero nonlinearity parameter $g$, gives rise to a plethora of interaction-induced phenomena that have no counterpart in the linear regime. The synergy between topology and interaction has ushered in new avenues for exploring topological transport and bulk-edge correspondence. A quintessential example is the nonlinear Thouless pumping [48]. In this context, nonlinearity plays a pivotal role in quantizing transport through the formation of solitons and spontaneous symmetry-breaking bifurcations [see Figs. 1(b2), (c2), and (d2)]. We remark that the discontinuous soliton position jumps in Figs. 1(c2) and 2(c2) arise from self-crossing band structures [see Figs. 1(c1) and 2(c1)], where adiabaticity breaks due to nonlinearity-induced band’s bifurcation at critical nonlinearity. For weak nonlinearity, the motion of the soliton becomes topologically quantized, as illustrated in Figs. 1(b2). Specifically, the bulk-edge correspondence of the nonlinear RM Hamiltonian (1) exhibits a quantized displacement by two lattice sites per cycle [see Fig. 1(b2)], which is directly linked to the topology of the underlying band structures [see Fig. 1(b1)]. In contrast, no such quantized pumping of solitons occurs when $g=0$, as shown in Fig. 1(a2). As the nonlinearity strength $g$ increases, self-intersecting bands emerge in the intermediate nonlinear regime with the appearance of the loop structure [88] as shown in Fig. 1(c1), leading to a pumping effect that persists but is no longer quantized, as observed in Fig. 1(c2). When the nonlinearity further intensifies into the strong nonlinear regime [see Fig. 1(d1)] and the self-trapped phenomenon is supposed to occur [89], the soliton dynamics are suppressed, resulting in the cessation of the pumping effect, as depicted in Fig. 1(d2)
@@
Figure 2: Eigenvalue’s nonlinearity of nonlinear RM Hamiltonian defined the last line of Table 1 and nonlinear Thouless pumping of soliton. (a1)-(d1): The auxiliary $\lambda$ spectrum for the nonlinear RM Hamiltonian (1). (a2)-(d2): The anticipated position of the nonlinear excitation of soliton as a function of time over one period. The parameters are fixed at $J=1$, $\delta=0.5$, $\Delta=1$, $T=2000\pi$ and $\omega=10^{-3}$. Specifically, (a1)-(d1) and their corresponding (a2)-(d2) panels represent interaction strengths of $g=0$, $g=1$, $g=5$, and $g=10$, respectively.
(iii) The eigenvalue’s nonlinearity in a linear Hamiltonian as summarized in the third line of Table 1: This scenario introduces a concept where the eigenvalues exhibit a state-variable-dependent nonlinearity. The eigenvalue equation is given by $\hat{H}_{\text{lin}}\Psi=\omega S(\omega)\Psi$. Here, the Hermitian matrix $H_{\text{lin}}$ and $S$ is the linear Hamiltonian and overlap matrix, respectively, $\Psi$ is the eigenvector. Note that the matrices $H_{\text{lin}}$ and $S$ may depend on the eigenvalue $\omega$, which is real. Remarkably, the authors of Ref. [64] showed that when the nonlinearity is weak but finite, the topological edge states of the auxiliary eigenstates are topologically inherited as physical edge states. This inheritance occurs when the auxiliary eigenvalues are monotonic with respect to the physical eigenvalues. This finding establishes a bulk-edge correspondence in systems with eigenvalue nonlinearity, highlighting the profound implications of such nonlinearity on the topological properties of the system.
(iv) The anomalous eigenvalue’s nonlinearity of nonlinear Hamiltonian in the fourth line of Table 1: This scenario extends the concept of eigenvalue nonlinearity from linear Hamiltonians in Ref. [64] to their nonlinear counterparts, which is referred to as the anomalous eigenvalue’s nonlinearity of the nonlinear Hamiltonian. The definition given by $\hat{H}_{\text{non}}\Psi=\omega S(\omega)\Psi$ represents a generalization where the Hamiltonian is nonlinear. Unlike systems with only wavefunction nonlinearity (e.g., $g|\Psi_{n}|^{2}$), the anomalous case involves dual nonlinearity, both the Hamiltonian ($H_{\text{non}}$) and eigenvalues [$\omega S(\omega)$] depend on the state. This equation models systems with anharmonic potentials, where energy levels and wave functions are intricately intertwined in a nonlinear fashion, providing a richer framework for understanding quantum states.